Minutes of the BEAST meeting
Friday January 22nd 12:00 JST

  • Schedule :

    At the last KEK-B steering committee, the commissioning run was extended until March 5. Further extensions of the run are possible but not yet decided. The next meeting on scheduling is February 15th. Constraints on the schedule include power consumption and the dedicated photon factory run.

  • KEK-B General:

    Currently, a major problem is the reproducibility of linac settings when switching from e+ to e-. Stable currents in both the HER and LER of over 10 mA have been achieved. The maximum current is limited by radiation safety. After the installation of Lambertson magnets for the abort system, this is no longer a constraint. At present, the emphasis is making collisions before the end of January. We should therefore prepare the EFC back to back trigger to be ready for collisions.

    For BEAST we have had one long coast for LER background measurement at a current of below 1 mA. A higher current LER fill is needed for a good first determination of the LER background.

    Closer cooperation between BEAST and KEK-B is needed in the future. Haba-san said that a joint workshop is planned around the week of February 15th.

  • Shifts and Data Taking:

    We will not record NTUPLEs during KEK-B machine experiments. The BEAST DAQ will run with run number=0 and the shifter still must be present to keep the DAQ alive, reset HV trips etc. Once stable fills are routine, we will record all BEAST data. We have requested help from the cosmic ray shift crew to man BEAST shifts in the coming months.

  • SVD Ladders in BEAST:

    The question discussed was whether SVD ladders in BEAST are needed to determine whether BELLE can be rolled in. Without SVD ladders, we will depend entirely on simulation to determine the SVD occupancy. Bozek noted that real beam is needed to fully test the sparsification algorithm. He noted that the DAQ/readout for the BEAST ladders can be considerably simpler than the DAQ used in the cosmic ray setup. Yamamoto discussed CLEO 2.5 problems that did not become apparent until after the SVX was installed with beam. This problems were related to timing in high occupancy events. Olsen noted that Peter Weilhammer had recommended that we get experience with real SVD ladders in beam as soon as possible. However, to install and readout ladders will require resources from the SVD group, a plan and a schedule. The BEAST group requests that this question be discussed and decided at the next executive board meeting.

  • BEAST Hardware Status:

    Diodes: Sahu reported on current hardware problems. During high background commissioning, 3 of the 24 PIN diodes were killed. One was lost due during HER commissioning when the tungsten collimator was in place. Two others were lost during the start of LER commissioning. During these high background tuning periods, a conductive path is established across the diode which makes it impossible to deplete the diode. To avoid this problem in the future, the 60 V depletion voltage has been reduced to 3 V until stable fills are achieved. The signals from the diodes will still be useful for accelerator tuning. (Recall the depletion depth is proportional to sqrt(V), so 1/4 of the wafer is depleted). In addition, Marc Rosen will prepare spare diode assemblies which can be installed during a future down time.


    MOSFETs: The integrated doses is now 20 krad from commissioning. The more sensitive R-type MOSFETs will soon run out of dynamic range. Two MOSFETs have been installed on the EFC. MOSFETs should also be installed on the CsI(Tl) crystals.

    Drift Tubes: The HV calibration has been done. The radiation dose calculated from the drift tube leakage current agrees with the diode dose to within a factor of two. To preserve the drift tubes, the high voltage has been reduced from the optimal 1.8-1.9 kV range to 1.6 kV. The trip current is currently been set to 10 microA. This can probably be increased to 100 microA.

    X-ray detector: A 32 channel pulse inverter was constructed. Amplifier and ADC were found. Now trying to get spectra.

    EFC: The scalers for the EFC are now understood. 2/8 scalers appear to be dead. This is under investigation.

  • DAQ Status:

    Moffit reported that no DAQ crashes have been reported during recent shifts. Bugs in NTUPLE filling have been corrected. Problems with out of range IOTECH channels are fixed. The BEAST display has been updated to take the lower diode voltage into account.

  • Simulation:

    Bozek reported on simulation work to understand the December 25th HER BEAST run. Taking into the account the vacuum pressure which was x 2 greater than previously assumed, and the 5 mm offset of the beam in the vertical direction, the simulation is within a factor of 2 of the observed Background/current/nTorr.

    However, he is unable to reproduce the angular distributions and front/back diode ratio with the GEANT simulation. Sahu's EGS simulation with lower energy cutoffs however reproduces the observed front/back ratio well. The effect of the vertical emittance (x 5 greater than nominal) on the background appears to be small. The effect of showering upstream of the nominal GEANT volume should be checked.